# Neutrino flavor instabilities in neutron star mergers with moment transport: slow, fast, and collisional modes

### v1

@@ Acknowledgements.
We thank Evan Grohs, Damiano Fiorillo and Tianshu Wang for valuable discussions, with special thanks to Tianshu Wang for generously providing the data used in Appendix C. J.F. acknowledges support from the Severo Ochoa Excellence Grant CEX2023-001292-S funded by MICIU/AEI/10.13039/501100011033, and from the Network for Neutrinos, Nuclear Astrophysics and Symmetries (N3AS), through the National Science Foundation Physics Frontier Center award No. PHY-2020275. J.P.K. and G.C.M. are supported by the United States Department of Energy, Office of Science, Office of Nuclear Physics under award numbers DE-FG02-02ER41216 and DE-SC00268442 (ENAF). G.C.M. also acknowledges support from DOE contracts LA22-ML-DE-FOA-2440 and DE-AC52-107NA27344. F.F. is supported by the Department of Energy, Office of Science, Office of Nuclear Physics, under contract number DE-SC0020435 and by NASA through grant 80NSSC22K0719.
@@ A.1 Eigenvalues of the stability matrix
With these assumptions, the now $8\times 8$ stability matrix is split into $4$ independent blocks. If we order the vector of perturbation amplitudes as $\left(\mathcal{A}_{ex},\bar{\mathcal{A}}_{xe},\mathcal{B}_{ex}^{\mathrm{x}},\overline{\mathcal{B}}_{xe}^{\mathrm{x}},\mathcal{B}_{ex}^{\mathrm{y}},\overline{\mathcal{B}}_{xe}^{\mathrm{y}},\mathcal{B}_{ex}^{\mathrm{z}},\overline{\mathcal{B}}_{xe}^{\mathrm{z}}\right)^{T}$, the matrix $\mathsf{S}$ has a block structure given by
@@
The subscripts “pres” and “break” refer to the “isotropy-preserving” and “isotropy-breaking” modes, following the nomenclature of [41, 43, 44]. For consistency with those works, we restrict to CFI and take $\omega=0$ in this section unless otherwise specified.
@@ A.2 Gapped/Minus and Gapless/Plus modes
As can be seen in Eqs. (43) and (45), each branch of the instability admits two modes, which have been dubbed in the literature “plus” and “minus” modes. Although it may be tempting to associate these modes to the signs ++ and - appearing in those equations, it would be erroneous since the identification of plus/minus modes depends on the sign of the various quantities (see for instance Appendix A in [67] [^8] We caution the reader that the expressions in [67] are for the physical frequency $\Omega$, not the shifted quantity $\Omega^{\prime}$, which differs by its real part.). A more physically-grounded nomenclature of these modes was used in Ref. [49], introducing the terms “gapless” and “gapped” modes.
@@
and since $\Omega^{\prime}_{+}\to 0$ for $\Gamma,\overline{\Gamma}\to 0$, it is a gapless mode. Note that the same classification holds if one includes the vacuum term, as the notion of gapless/gapped depends upon the limit $\Gamma,\overline{\Gamma}\to 0$ and $\Delta m^{2}\to 0$ [49].
@@ B.3 Comparison of the two methods
For a general polychromatic spectrum, it is possible to solve numerically Eq. (55) for $\Omega^{\prime}$, or more specifically the discretized version (52)–(54). For gapped modes, each term of the sum is of order 1, since $\mathrm{Re}(\Omega^{\prime})\simeq\sum_{n}(\overline{\Delta}_{N}^{(n)}-\Delta_{N}^{(n)})\Delta p_{n}$ [see Eq. (43), valid for a gapped mode]. This is easy to handle numerically. However, for gapless modes, the denominator of each term is typically of order $\Gamma$ (since the real part of $\Omega^{\prime}$ is very small compared to the imaginary part, itself of order $\sim\Gamma$). This means that each term in the sum is roughly of order $\sqrt{2}G_{F}\mathcal{N}/\Gamma\sim 10^{4}$ and above. Numerically, dealing with a sum of very large terms that must be precisely combined to give 1-1 or $3$ is more difficult. As such, we have found that in a few cases, it was important to provide a good seed to the root solver to avoid missing a gapless instability when using Eq. (55). Apart from this precaution, finding a root of the integral equation or using the stability matrix method are both reliable methods.
Numerically solving Eq. (55) using the routine scipy.optimize.root [118] for a single point takes $\sim$ 1 ms of CPU time (including both the 1-1 and $3$ branches). However, since at each point, Eq. (55) can have multiple solutions for $\Omega^{\prime}$, we can use multiple initial (seed) guesses to capture all the solutions (see also the comment above for possible difficulties with gapless modes). We can use an ad hoc set of 10 purely imaginary seed guesses $\mathrm{i}\left[10^{0},10^{1},\dots,10^{9}\right]\mathrm{s}^{-1}$, at equal logarithmic intervals, that covers the typical range of any growth rate that might occur. This multiple seed guesses method scales up the computation time by one order of magnitude. The diagonalization of the stability matrix (with the routine scipy.linalg.eig [118]) for a single point takes $\sim 0.2\,\mathrm{ms}$ if we restrict to the isotropic case [that is, for the multi-energy version of (41)], and $\sim 8\,\mathrm{ms}$ for the full stability matrix. The stability matrix method is thus very competitive with the dispersion relation approach, since it includes nonzero fluxes, doesn’t require any guesses and is as reliable for gapped and gapless modes. All computation times discussed here were for 16 energy groups, that is the energy binning of the NuLib table introduced in Sec. III.1.
Figure 14: Comparison of typical CFI gapless modes in CCSN and NSM environments. The left panels show the energy spectra and collision rates for the CCSN example, which corresponds to Fig. 1 in [47]. The middle panels show the same quantities for the point identified with an orange star on Fig. 2. On the right panel, we show the result of multi-energy and energy-averaged LSA for these two configurations.
@@
To highlight the difference between those two astrophysical environments, we focus on two simulation points. The first one is studied in Fig. 1 of Ref. [47], and corresponds to the configuration at a radius of 37 km, at 120 ms post-bounce for a $18M_{\odot}$ 1D CCSN simulation. The second point is shown with an orange star in Fig. 2, it has coordinates $(X,Y,Z)\simeq(-33\,\mathrm{km},0\,\mathrm{km},-5\,\mathrm{km})$. It is chosen because it is in the middle of a gapless region that remains a CFI region even when including anisotropies (see Fig. 11). The energy spectra of (anti)neutrinos and the collision rates are shown on the left and middle panels of Fig. 14. On the right panel, we report the values of $\mathrm{Im}(\Omega)_{\mathrm{max}}$ obtained assuming homogeneity and isotropy, for the multi-energy and the two energy-averaged approaches discussed in Sec. IV.1.
@@
A key difference with the 7 ms postmerger snapshot is that in the 3 ms snapshot there is a system-wide overabundance of $\bar{\nu}_{e}$ over $\nu_{e}$, which is consistent with an early phase of protonization. At 7 ms, the system is closer to a slow, quasi-steady-state evolution where some regions are protonizing and others are neutronizing. A consequence of $\overline{\mathcal{N}}_{ee}>\mathcal{N}_{ee}$ is that only gapped modes can occur, and we know they can be described correctly with the monochromatic method A. Nevertheless, we see in Fig. 15 that the CFI is mostly subdominant, even in the anisotropic case (middle panel), where anisotropy-driven CFIs appear at large $|Z|$, but with very small growth rates. The inclusion of the vacuum term doesn’t dramatically alter the picture, even though they lead to larger growth rates than the CFI at larger distances.
Figure 15: Flavor instabilities in the 3 ms postmerger snapshot from the same NSM simulation studied in the main text. Top: results assuming isotropic distributions and without a vacuum term, showing the gapless (orange tones) and gapped (blue tones) CFI regions. Middle: results from moment LSA at $\mathbf{k}^{\prime}=\mathbf{0}$, without a vacuum term. Bottom: same as middle panel but including the vacuum term (NO), and superimposing the FFI estimate (38).
@@ D.2 Different component masses
Finally, we consider a 5 ms postmerger snapshot from an independent M1 simulation, taken from Ref. [92]. This simulation of a neutron star merger with component masses of $1.2M_{\odot}$ and $1.2M_{\odot}$ was performed with an earlier version of the two-moment radiation transport code SpEC also used in [52]. While in the simulation studied in the main text, the HMNS collapses into a black hole at $t\sim 8.5\,\mathrm{ms}$, the simulation in [92] shows no sign of collapse in its $10\,\mathrm{ms}$ of evolution. However, this snapshot is very similar to the 7 ms one studied in the main text, with thin bands of mostly gapless CFI near the surface of the HMNS, and alternance of gapless and gapped CFI regions in the tidal arms. Although we do not show it here, the qualitative and quantitative agreement of the monochromatic CFI estimate (35) is excellent, with once again a ratio of gapless growth rates of about $\sim 2$ between the method B and multi-energy estimates.
@@
Similarly to what we see in Fig. 12, when including anisotropies the CFI extends the instability regions in the disk, while slow modes appear in the polar regions. There are also likely slow mode regions at large $|X|$, but the simulation data available does not extend to that region. The instability landscape discussed in this work thus seems to generally extend to “late” postmerger moment-based simulations, although using snapshots from completely independent simulation codes would be needed to definitely confirm that trend.

### v2

@@ Acknowledgements.
We thank Evan Grohs, Damiano Fiorillo, and Tianshu Wang for valuable discussions, with special thanks to Tianshu Wang for generously providing the data used in Appendix C. J.F. acknowledges support from the Severo Ochoa Excellence Grant CEX2023-001292-S funded by MICIU/AEI/10.13039/501100011033 and from the Network for Neutrinos, Nuclear Astrophysics and Symmetries (N3AS), through the National Science Foundation Physics Frontier Center Award No. PHY-2020275. J.P.K. and G.C.M. are supported by the United States Department of Energy, Office of Science, Office of Nuclear Physics under Awardx No. DE-FG02-02ER41216 and No. DE-SC0024388 (ENAF). G.C.M. also acknowledges support from DOE Contracts No. LA22-ML-DE-FOA-2440 and No. DE-AC52-107NA27344. F.F. is supported by the Department of Energy, Office of Science, Office of Nuclear Physics, under Contract No. DE-SC0020435 and by NASA through Grant No. 80NSSC22K0719.
@@ A.1 Eigenvalues of the stability matrix
With these assumptions, the now $8\times 8$ stability matrix is split into four independent blocks. If we order the vector of perturbation amplitudes as $\left(\mathcal{A}_{ex},\bar{\mathcal{A}}_{xe},\mathcal{B}_{ex}^{\mathrm{x}},\overline{\mathcal{B}}_{xe}^{\mathrm{x}},\mathcal{B}_{ex}^{\mathrm{y}},\overline{\mathcal{B}}_{xe}^{\mathrm{y}},\mathcal{B}_{ex}^{\mathrm{z}},\overline{\mathcal{B}}_{xe}^{\mathrm{z}}\right)^{T}$, the matrix $\mathsf{S}$ has a block structure given by
@@
The subscripts “pres” and “break” refer to the “isotropy-preserving” and “isotropy-breaking” modes, respectively, following the nomenclature of [41, 43, 44]. For consistency with those works, we restrict to CFI and take $\omega=0$ in this section unless otherwise specified.
@@ A.2 Gapped/minus and gapless/plus modes
As can be seen in Eqs. (43) and (45), each branch of the instability admits two modes, which have been dubbed in the literature “plus” and “minus” modes. Although it may be tempting to associate these modes to the signs ++ and - appearing in those equations, it would be erroneous since the identification of plus and minus modes depends on the sign of the various quantities (see [^8] We caution the reader that the expressions in [68] are for the physical frequency $\Omega$, not the shifted quantity $\Omega^{\prime}$, which differs by its real part. for instance Appendix A in [68]). A more physically grounded nomenclature of these modes was used in Ref. [49], introducing the terms “gapless” and “gapped” modes.
@@
and since $\Omega^{\prime}_{+}\to 0$ for $\Gamma,\overline{\Gamma}\to 0$, it is a gapless mode. Note that the same classification holds if one includes the vacuum term, as the notion of gapless and gapped depends upon the limit $\Gamma,\overline{\Gamma}\to 0$ and $\Delta m^{2}\to 0$ [49].
@@ B.3 Comparison of the two methods
For a general polychromatic spectrum, it is possible to solve numerically Eq. (55) for $\Omega^{\prime}$ or, more specifically, the discretized version (52)–(54). For gapped modes, each term of the sum is of order 1, since $\mathrm{Re}(\Omega^{\prime})\simeq\sum_{n}(\overline{\Delta}_{N}^{(n)}-\Delta_{N}^{(n)})\Delta p_{n}$ [see Eq. (43), valid for a gapped mode]. This is easy to handle numerically. However, for gapless modes, the denominator of each term is typically of order $\Gamma$ (since the real part of $\Omega^{\prime}$ is very small compared to the imaginary part, itself of order $\sim\Gamma$). This means that each term in the sum is roughly of order $\sqrt{2}G_{F}\mathcal{N}/\Gamma\sim 10^{4}$ and above. Numerically, dealing with a sum of very large terms that must be precisely combined to give 1-1 or $3$ is more difficult. As such, we have found that in a few cases, it was important to provide a good seed to the root solver to avoid missing a gapless instability when using Eq. (55). Apart from this precaution, finding a root of the integral equation or using the stability matrix method are both reliable methods.
Numerically solving Eq. (55) using the routine scipy.optimize.root [129] for a single point takes $\sim$ 1 ms of CPU time (including both the 1-1 and $3$ branches). However, since at each point Eq. (55) can have multiple solutions for $\Omega^{\prime}$, we can use multiple initial (seed) guesses to capture all the solutions (see also the comment above for possible difficulties with gapless modes). We can use an ad hoc set of 10 purely imaginary seed guesses $\mathrm{i}\left[10^{0},10^{1},\dots,10^{9}\right]\mathrm{s}^{-1}$, at equal logarithmic intervals, that covers the typical range of any growth rate that might occur. This multiple seed guesses method scales up the computation time by one order of magnitude. The diagonalization of the stability matrix (with the routine scipy.linalg.eig [129]) for a single point takes $\sim 0.2\,\mathrm{ms}$ if we restrict to the isotropic case [that is, for the multi-energy version of (41)], and $\sim 8\,\mathrm{ms}$ for the full stability matrix. The stability matrix method is thus very competitive with the dispersion relation approach, since it includes nonzero fluxes, does not require any guesses and is as reliable for gapped and gapless modes. All computation times discussed here were for 16 energy groups, that is the energy binning of the NuLib table introduced in Sec. III.1.
Figure 14: Comparison of typical CFI gapless modes in CCSN and NSM environments. The left panels show the energy spectra and collision rates for the CCSN example, which corresponds to Fig. 1 in [47]. The middle panels show the same quantities for the point identified with an orange star in Fig. 2. In the right panel, we show the result of multi-energy and energy-averaged LSA for these two configurations.
@@
To highlight the difference between those two astrophysical environments, we focus on two simulation points. The first one is studied in Fig. 1 of Ref. [47] and corresponds to the configuration at a radius of 37 km, at 120 ms post-bounce for a $18M_{\odot}$ 1D CCSN simulation. The second point is shown with an orange star in Fig. 2, it has coordinates $(X,Y,Z)\simeq(-33\,\mathrm{km},0\,\mathrm{km},-5\,\mathrm{km})$. It is chosen because it is in the middle of a gapless region that remains a CFI region even when including anisotropies (see Fig. 11). The energy spectra of (anti)neutrinos and the collision rates are shown in the left and middle panels of Fig. 14. In the right panel, we report the values of $\mathrm{Im}(\Omega)_{\mathrm{max}}$ obtained assuming homogeneity and isotropy, for the multi-energy and the two energy-averaged approaches discussed in Sec. IV.1.
@@
A key difference with the 7 ms postmerger snapshot is that in the 3 ms snapshot there is a system-wide overabundance of $\bar{\nu}_{e}$ over $\nu_{e}$, which is consistent with an early phase of protonization. At 7 ms, the system is closer to a slow, quasi-steady-state evolution where some regions are protonizing and others are neutronizing. A consequence of $\overline{\mathcal{N}}_{ee}>\mathcal{N}_{ee}$ is that only gapped modes can occur, and we know they can be described correctly with the monochromatic method A. Nevertheless, we see in Fig. 15 that the CFI is mostly subdominant, even in the anisotropic case (middle panel), where anisotropy-driven CFIs appear at large $|Z|$, but with very small growth rates. The inclusion of the vacuum term does not dramatically alter the picture, even though they lead to larger growth rates than the CFI at larger distances.
Figure 15: Flavor instabilities in the 3 ms postmerger snapshot from the same NSM simulation studied in the main text. Top: results assuming isotropic distributions and without a vacuum term, showing the gapless (orange tones) and gapped (blue tones) CFI regions. Middle: results from moment LSA at $\mathbf{k}^{\prime}=\mathbf{0}$, without a vacuum term. Bottom: the same as the middle panel but including the vacuum term (NO), and superimposing the FFI estimate (38).
@@ D.2 Different component masses
Finally, we consider in Fig. 16 a 5 ms postmerger snapshot from an independent M1 simulation, taken from Ref. [93]. This simulation of a neutron star merger with component masses of $1.2M_{\odot}$ and $1.2M_{\odot}$ was performed with an earlier version of the two-moment radiation transport code SpEC also used in [52]. While in the simulation studied in the main text, the HMNS collapses into a black hole at $t\sim 8.5\,\mathrm{ms}$, the simulation in [93] shows no sign of collapse in its $10\,\mathrm{ms}$ of evolution. However, this snapshot is very similar to the 7 ms one studied in the main text, with thin bands of mostly gapless CFI near the surface of the HMNS, and alternance of gapless and gapped CFI regions in the tidal arms. Although we do not show it here, the qualitative and quantitative agreement of the monochromatic CFI estimate (35) is excellent, with once again a ratio of gapless growth rates of about $\sim 2$ between the method B and multi-energy estimates.
@@
Similarly to what we see in Fig. 12, when including anisotropies the CFI extends the instability regions in the disk, while slow modes appear in the polar regions. There are also likely slow mode regions at large $|X|$, but the simulation data available do not extend to that region. The instability landscape discussed in this work thus seems to generally extend to “late” postmerger moment-based simulations, although using snapshots from completely independent simulation codes would be needed to definitely confirm that trend.